# On the Incompressibility of Truth With Application to Circuit Complexity

### v1

@@ 1 Introduction
In this paper, we revisit the fundamentals of Circuit Complexity and the nature of efficient computation from a new perspective. We present a framework for understanding Circuit Complexity through the lens of Information Theory with analogies to results in Kolmogorov Complexity, viewing circuits as descriptions of truth tables, encoded in logical gates and wires, rather than purely computational devices. From this framework, we re-prove some existing strong Circuit Complexity bounds, explain what the optimal circuits for most boolean functions look like structurally, give insight into new circuit bounds, and explain the aforementioned results in a unifying intuition that re-frames time entirely.
@@ 1.1 Brief Classical View of Circuits
1.2 New View of Circuits as Descriptions
@@ 1.2.3 Connection to Kolmogorov Complexity
The Kolmogorov complexity [2] $K(s)$ of a string $s$ is the length of the shortest program (in some fixed universal language) that outputs $s$.
@@ 2 Applying the framework to existing bounds
Having established circuits as logical descriptions of truth tables, we now revisit classical results through this new lens. Rather than prove new theorems yet, we first demonstrate that known results emerge naturally, and with potentially greater clarity from the information-theoretic perspective.
This serves two purposes: validates that the framework is mathematically sound, capable of recovering established facts as special cases, and it reveals why these results hold, not merely that they hold. The explanations we obtain are potentially more intuitive than the original proofs, suggesting that the descriptive view is not merely equivalent to the computational view, but is potentially more fundamental. We begin with the cornerstone result of circuit complexity: Shannon’s 1949 lower bound [6].
@@ 2.1 Shannon’s Lower Bound through Information Theory
On the other hand, by the basic entropy bound for lossless codes (or by Shannon’s source coding theorem [5], using a prefix-free refinement of our encoding) any lossless description scheme must satisfy
@@
$s\log s\;\geq\;c^{\prime}\,2^{n}\quad\Rightarrow\quad s\;\geq\;\Omega\!\Big(\frac{2^{n}}{\log(2^{n})}\Big)\;=\;\Omega\!\Big(\frac{2^{n}}{n}\Big).$
@@
Most truth tables are essentially “random”: they have barely any structure for logic to compress. Random data is mostly incompressible. Therefore, the circuit “description” must be nearly as large as the data itself. The factor of $n$ in the denominator ($2^{n}/n$ rather than $2^{n}$) represents the minimal overhead of circuit structure: the “syntax cost” of writing gates and wires instead of raw bits. It’s the best you can do even with optimal organization. This is not a non-constructive combinatorial coincidence. It is an information-theoretic necessity.
@@ 2.2 What Do Optimal Circuits Look Like?
Having established that most boolean functions need $\Omega(2^{n}/n)$ gates, we can now ask: what is the structure of these optimal circuits? Answering this question seems mysterious from a purely combinatorial view. However, the information-theoretic view provides an intuitive prediction: if a circuit cannot compress the truth table, it must essentially enumerate it, essentially become the table itself.
@@
This prediction is not new: it was already implicitly known through Lupanov’s representation [3]. However, the framework reveals why this structure is optimal: it’s similar to the most efficient way to “print” a truth table using logical gates.
@@
The construction in the canonical form naively uses $O(2^{n})$ gates (one detector per true row, each with $n$ gates). Lupanov’s optimization organizes this like a decision tree: check variables sequentially, reusing intermediate comparisons across branches. For brevity and sake of not reinventing, we will not go into great detail of the construction of the representation, but cover enough to connect to the intuition.
@@
The prediction confirmed: Shannon’s and Lupanov’s results together say that, for the very hardest functions, optimal circuits are essentially optimized lookup tables: compressed representations that approach the theoretical limit of incompressibility. In particular, many circuits admit isomorphic rewritings, so even if we had an algorithm for computing minimum circuits for a given truth table, the circuit it returned could be viewed as an “encryption” of the function’s true nature, hiding the fact that it is logically equivalent to a short-circuited LUT. A natural notion of “decryption” would be to test whether the circuit is isomorphic to an optimized lookup-table structure. In this sense, what looks like clever algebraic manipulation might be much simpler in description: it is just the best possible way to wire the truth table into gates. In this view, non-uniform efficient computability is not merely related to description length — it is defined by it: a function is efficiently computable in the circuit model precisely to the extent that its truth can be compressed into a small circuit. Time, in this picture, shows up only as a secondary constraint on how those compressed descriptions are discovered or used, not on what is in principle computable.
@@ 2.3 Revisiting the Natural Proofs Barrier
In 1993, Razborov and Rudich [4] analyzed successful circuit lower bound techniques from the 1980s and showed that if these techniques were used to prove separations like $P\neq NP$, then we could use the properties of those proofs to obtain faster distinguishers of random functions from pseudo-random functions.
@@ 2.4 Revisiting Proof Complexity Struggles
In essence, if there is a polynomial-bounded proof system that proves, for every $T$, that “$T$ is incompressible,” then that proof system looks to be compressing the incompressible. Imagine a hypothetical polynomial SAT solver somehow using a concise logical description to describe every incompressible logical description on $n$ variables.
@@ 3 Investigating MCSP
Having seen how the information-theoretic framework illuminates classical results, we now examine the Minimum Circuit Size Problem (MCSP) that has long had a mysterious complexity. The decision version of this problem asks, given a truth table, $T$, and a size bound, $s$, whether the table can be computed by a circuit of at most $s$ gates. MCSP is a meta-problem that has been studied extensively for its connections to cryptography, learning theory, and derandomization. Various conditional hardness results exist, but unconditional lower bounds have been elusive.
@@
MCSP is in $NP$ (the class of problems where solutions are verifiable in polynomial time in the size of its input [1]).
@@ 3.2 The Information-Theoretic Impossibility
  • Slightly compressible: Circuit sizes like $2^{n/k}$ for moderate $k$. Some structure, but exponential complexity. An intermediate regime.
@@
Suppose, for contradiction, that a circuit $C$ of size $poly(2^{n})$ (polynomial in the truth table size) solves $MCSP_{SEARCH}$. Then $C$ is effectively a short description that encodes the shortest description of every table ($2^{2^{n}}$ of them) on $n$ variables, where the descriptions are in the “circuit compression scheme.”
This is like asking $C$ to “compress the incompressible” as every hard table’s information on $n$ variables is encoded in $C$.
@@
$O(poly(2^{n})\cdot\log(poly(2^{n}))\ll\Omega(\frac{2^{c\cdot 2^{n}}}{2^{n}}\cdot\log(\frac{2^{c\cdot 2^{n}}}{2^{n}}))$ where $1>c>0$ and represents an “entropy fence” the circuit can’t pass.
@@
With a relatively “small” $MCSP_{SEARCH}$ circuit, we could input any data we’d like as bits (a truth table, a book, DNA, etc) and it would efficiently output the most efficient way to store it in the circuit logic scheme. This makes $MCSP_{SEARCH}$’s efficiency not only worthwhile for problem solving, but also for perfectly optimal lossless information storage. An efficient $MCSP_{SEARCH}$ circuit is an efficient universal data compressor, which sounds like a fantasy.
@@
Construct the table of the function, $f:\{0,1\}^{2^{n}}\rightarrow Enc(C_{n})$, that defines the optimal circuit compressor, where for each row, $Enc(C_{N})$ is the bit encoding of the optimal circuit description of the $\{0,1\}^{2^{n}}$ bits on the left-side of the row for a fixed basis, such as $\{AND,OR,NOT\}$.
@@
This is feasible for small $n$ and we can use modern circuit minimization algorithms to find the smallest circuit computing/describing $f$. If the minimal circuits obtained look more like an exhaustive description of every pattern in the tables rather than an algorithm, the hypothesis survives for small $n$.
@@
We can tap into what the “Book of Books” looks like on a small scale.

### v4

@@ 1 Introduction
In this paper, we revisit the fundamentals of Circuit Complexity and the nature of efficient computation from a fresh perspective. We present a framework for understanding Circuit Complexity through the lens of Information Theory with analogies to results in Kolmogorov Complexity, viewing circuits as descriptions of truth tables, encoded in logical gates and wires, rather than purely computational devices. From this framework, we re-prove some existing Circuit Complexity bounds, explain what the optimal circuits for most boolean functions look like structurally, give an explicit boolean function family that requires exponential circuits, and explain the aforementioned results in a unifying intuition that re-frames time entirely.
@@ 1.1 Brief Classical View of Circuits
1.2 View of Circuits as Descriptions
@@ 1.2.3 Connection to Kolmogorov Complexity
The Kolmogorov complexity [undefb] $K(s)$ of a string $s$ is the length of the shortest program (in some fixed universal language) that outputs $s$.
@@ 2 Applying the framework to existing bounds
Having established circuits as logical descriptions of truth tables, we now revisit classical results through this lens. Rather than prove new theorems yet, we first demonstrate that known results emerge naturally, and with potentially greater clarity from the information-theoretic perspective.
This serves two purposes: validates that the framework is mathematically sound, capable of recovering established facts as special cases, and it reveals why these results hold, not merely that they hold. The explanations we obtain are potentially more intuitive than the original proofs, suggesting that the descriptive view is not merely equivalent to the computational view, but is potentially more fundamental. We begin with the cornerstone result of circuit complexity: Shannon’s 1949 lower bound [undeff].
@@ 2.1 Shannon’s Lower Bound through Information Theory
On the other hand, by the basic entropy bound for lossless codes (or by Shannon’s source coding theorem [undefe], using a prefix-free refinement of our encoding) any lossless description scheme must satisfy
@@
$s\log s\;\geq\;c\,2^{n}\quad\Rightarrow\quad s\;\geq\;\Omega\!\Big(\frac{2^{n}}{\log(2^{n})}\Big)\;=\;\Omega\!\Big(\frac{2^{n}}{n}\Big).$
@@
Most truth tables are essentially “random”: they have barely any structure for logic to compress. Random data is mostly incompressible. Therefore, the circuit “description” must be nearly as large as the data itself. The factor of $n$ in the denominator ($2^{n}/n$ rather than $2^{n}$) represents the minimal overhead of circuit structure: the “syntax cost” of writing gates and wires instead of raw bits. It’s the best you can do even with optimal organization. This is not a non-constructive combinatorial coincidence. It is an information-theoretic necessity that shows logic cannot cheat entropy. Circuits store data, and this is even clearer in the context of Solid-State Drives (SSDs) that use $NAND$ flash memory.
@@ 2.2 What Do Optimal Circuits Look Like?
Having established that most boolean functions need $\Omega(2^{n}/n)$ gates, we can now ask: what is the structure of these optimal circuits? Answering this question seems mysterious from a purely combinatorial view. However, the information-theoretic view provides an intuitive prediction: if a circuit cannot compress the truth table, it must essentially become the table itself.
@@
This prediction is not new: it was already implicitly known through Lupanov’s representation [undefc]. However, the framework reveals why this structure is optimal: it’s similar to the most efficient way to “print” a truth table using logical gates.
@@
The construction in the canonical form naively uses $O(n\cdot 2^{n})$ gates (one detector per true row, each with $n$ gates). Lupanov’s optimization organizes this like a decision tree: check variables sequentially, reusing intermediate comparisons across branches. For brevity and sake of not reinventing, we will not go into great detail of the construction of the representation, but cover enough to connect to the intuition.
@@
The prediction confirmed: Shannon’s and Lupanov’s results together say that, for the very hardest functions, optimal circuits are essentially optimized lookup tables: compressed representations that approach the theoretical limit of incompressibility. The analogy to SSDs is even more interesting in the case of Lupanov’s representation because it uses “address blocks” and “payload blocks.” In this sense, what looks like clever algebraic manipulation might be much simpler in description: it is just the best possible way to wire the truth table into gates. In this view, non-uniform efficient computability is not merely related to description length — it is defined by it: a function is efficiently computable in the circuit model precisely to the extent that its truth can be compressed into a small circuit. Time, in this picture, shows up only as a secondary constraint on how those compressed descriptions are discovered or used, not on what is in principle computable.
@@ 2.3 Revisiting the Natural Proofs Barrier
In 1993, Razborov and Rudich [undefd] analyzed successful circuit lower bound techniques from the 1980s and showed that if these techniques were used to prove separations like $P\neq NP$, then we could use the properties of those proofs to obtain faster distinguishers of random functions from pseudo-random functions.
@@ 2.4 Revisiting Proof Complexity Struggles
In essence, if there is a polynomial-bounded proof system that proves, for every $T$, that “$T$ is incompressible,” then that proof system looks to be compressing the incompressible. In Section 4, we provide functions that are maximally hard for circuits, and so they are likely good candidates for hard $MCSP$ tautologies/contradictions outlined above.
@@ 3 Investigating MCSP
Having seen how the information-theoretic framework illuminates classical results, we now examine the Minimum Circuit Size Problem (MCSP) that has long had a mysterious complexity. The decision version of this problem asks, given a truth table, $T$, and a size bound, $s$, whether the table can be computed by a circuit of at most $s$ gates. MCSP is a meta-problem that has been studied extensively for its connections to cryptography, learning theory, and derandomization.
@@
MCSP is in $NP$ (the class of problems where solutions are verifiable in polynomial time in the size of its input [undefa]).
@@ 3.2 The Information-Theoretic Impossibility
  • Slightly compressible: Circuit sizes like $2^{n/k}$ for moderate $k$. Some structure, but exponential complexity. An intermediate regime that are likely more condensed lookup tables.
@@
If a circuit $C$ of size $poly(2^{n})$ solves $MCSP_{SEARCH}$, then $C$ is effectively a short circuit description that encodes the shortest circuit description of every table ($2^{2^{n}}$ of them) on $n$ variables.” This is like asking $C$ to “compress the incompressible.”
@@
$O(poly(2^{n})\cdot\log(poly(2^{n}))\ll\Omega(\frac{2^{c\cdot 2^{n}}}{2^{n}}\cdot\log(\frac{2^{c\cdot 2^{n}}}{2^{n}}))$ where $\frac{1}{2}>c>0$ and represents a sensitive “entropy fence” the circuit can’t pass.
@@
With a relatively “small” $MCSP_{SEARCH}$ circuit, we could input any data we’d like as bits (a truth table, a book, DNA, etc) and it would efficiently output the most efficient way to store it in the circuit logic scheme. This makes $MCSP_{SEARCH}$’s efficiency not only worthwhile for problem solving, but also for essentially optimal lossless information storage. An efficient $MCSP_{SEARCH}$ circuit is an efficient universal data compressor, which sounds like a fantasy.
@@
Construct the table of the function, $f:\{0,1\}^{2^{n}}\rightarrow Enc(C_{n})$, where for each row, $Enc(C_{N})$ is the bit encoding of the optimal circuit description of the $\{0,1\}^{2^{n}}$ bits on the left-side of the row for a fixed basis.
@@
If the minimal circuits obtained for $f$ look like an exhaustive description of every pattern in the tables, forced to pay an entropy bill rather than be a concise algorithm, the hypothesis survives for small $n$.
@@
We can tap into what the “Book of Summaries” looks like on a small scale.
4 An Explicit Boolean Function Family That Requires Exponential Circuits
We can now use the fact that circuits operate as lossless codes to prove an exponential lower bound on an explicit family of Boolean functions. To do this, we construct the Boolean function family, $\{f_{\Omega_{n}}\}$, where each $f_{\Omega_{n}}$ in the set has its output bits equal to the first $2^{n}$ bits of a Chaitin constant $\Omega_{U}$. A circuit for this problem decides, given a position of a bit (represented as a row in the truth table) in $\Omega_{U}$, whether the bit is 0 or 1. Essentially, it solves a finite version of the halting problem. Since the exact bits of $\Omega_{U}$ depend on the universal prefix-free Turing machine $U$, the $\{f_{\Omega_{n}}\}$ family is infinite.
Definition 4.0.1.
A Chaitin constant  $\Omega_{U}$ is the halting probability of a universal prefix-free Turing machine $U$ [undef], and it is algorithmically random, so the shortest program to output the first $n$ bits of  $\Omega_{U}$ must be of size at least nO(1)n-O(1).
Theorem 4.1.
The boolean function family, $\{f_{\Omega_{n}}\}$, requires 2-fan-in circuits of $\Omega(2^{n}/n)$ gates.
Proof.
Assume, for the sake of contradiction, for each $n$, there exists a circuit $C_{n}$ with $O(s)$ gates computing/describing $f_{\Omega_{n}}$. Then, we can store each $C_{n}$ with $O(s\log s)$ bits. Then, create a program, $P_{n}$, that contains the encoding of $C_{n}$ and decodes it into $C_{n}$ and decompresses $C_{n}$ into the first $2^{n}$ bits of  $\Omega_{U}$, and prints them (essentially what the first part of a verifier program for $MCSP$ does, but hardcoded). This means $P_{n}$ produces the first $2^{n}$ bits of  $\Omega_{U}$ as output with $|P_{n}|=O(s\log s)+O(1)$.
For sufficiently large $n$, $O(s\log s)\ll\Omega(2^{n})$ bits, so circuits encoded in $O(s\log s)$ bits for $\{f_{\Omega_{n}}\}$ would let us compress incompressible truth: a contradiction. Thus, the only way out is $s=\Omega(2^{n}/n)$. ∎
5 Additional Questions to Explore
Based on the results and conjectures above, the following problems would be interesting.
  • Define $\Omega_{U}^{t}$ as the probability that a universal prefix-free Turing machine $U$ halts within $t$ steps. If we know $\Omega_{U}^{t}$, after enough programs for $U$ have halted within $t$ steps, we know the rest do not, so $\Omega_{U}^{t}$ is like a bounded oracle, and its bits should be very pseudo-random.
    How compressible is $\Omega_{U}^{t}$ in circuit form?
    As $t$ approaches $\infty$, $\Omega_{U}^{t}$ becomes closer to $\Omega_{U}=\Omega_{U}^{\infty}$.
    We can also consider how the truth tables of bounded-time halting problems can be formulated to encode $\Omega_{U}^{t}$:
    • Recall the classic P-Complete and EXPTIME-Complete problems:
        $BH_{UNARY}=\{\langle M,x,1^{t}\rangle:M\textnormal{ accepts }x\textnormal{ within }t\textnormal{ steps}\}$
        $BH_{BINARY}=\{\langle M,x,t\rangle:M\textnormal{ accepts }x\textnormal{ within }t\textnormal{ steps}\}$ where $t$ is encoded in binary.
    • Now consider alternative formulations of them where $U_{0}$ is fixed and push $y:=\langle M,x\rangle$ as an encoding into the input string:
        $UBH_{UNARY}(U_{0})=\{\langle y,1^{t}\rangle:U_{0}(y)\textnormal{ accepts within }t\textnormal{ steps}\}$
        $UBH_{BINARY}(U_{0})=\{\langle y,t\rangle:U_{0}(y)\textnormal{ accepts within }t\textnormal{ steps}\}$ where $t$ is encoded in binary.
        Since $U_{0}$ can polynomially simulate any $M$, these fixed-machine formulations are still complete respectively.
    • Observe that $\Omega_{U}^{t}$ is literally a weighted aggregate of bounded-halting answers. Fix a specific prefix-free universal machine $U$ and a step model where in $t$ steps it can inspect at most $t$ input bits. Then:
        $\Omega_{U}^{t}=\textnormal{Pr}[U\textnormal{ halts within }t]=\sum\limits_{p:\,U(p)\downarrow\leq t}2^{-|p|}$
        Due to the “can’t read more than $t$ bits in $t$ steps” fact, any program that halts within $t$ must satisfy $|p|\leq t$. So define the finite truth table slice:
        $H_{t}(p):=1[U(p)\textnormal{ halts within }t],p\in\{0,1\}^{\leq t}$
        Then,
        $\Omega_{U}^{t}=\sum\limits_{p\in\{0,1\}^{\leq t}}2^{-|p|}H_{t}(p)$
        So in the strongest sense, the truth table of a bounded-halting problem contains $\Omega_{U}^{t}$ as a simple function of that truth table slice, suggesting $\Omega_{U}$’s entropy could trickle down as $\Omega(f(n))$ circuit lower bounds for its time-bounded forms because $\Omega_{U}^{t}$ should look sufficiently random to particular $t$-bounded truth compressors with $U$ acting like a seed to the pseudo-randomness.
  • Is it possible that, assuming non-uniform or uniform $poly(2^{n})$ circuits for $MCSP_{SEARCH}$ exist, we can extract sub-circuits to create small programs that output incompressible bits?
  • Can efficient Circuit Frege proofs for “incompressible” $MCSP$ tautologies get unrolled to create small programs that output incompressible bits?